Using the following chemical equation how many liters of carbon dioxide can be produced by burning 79 L of oxygen?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-16T07:41:48-0500

C4H10 + 5O2 -> 3CO2 + 4H2O - unbalanced equation


Solution:

The balanced chemical equation:

2C4H10 + 13O2 → 8CO2 + 10H2O

According to the equation above: n(O2)/13 = n(CO2)/8


At STP, one mole of any gas occupies a volume of 22.4 L (Vm).

Hence,

V(O2) / (13 × Vm) = V(CO2) / (8 × Vm)

V(O2) / 13 = V(CO2) / 8

To find the volume of carbon dioxide, solve the equation for V(CO2):

V(CO2) = 8 × V(O2) / 13 = 8 × 79 L / 13 = 48.615 L = 48.62 L

V(CO2) = 48.62 L


Answer: 48.62 liters of carbon dioxide (CO2) can be produced.


However, if an unbalanced equation is used, 47.4 liters of carbon dioxide (CO2) can be produced.